Determine the overall cell potential for a system in which the following equation represent s the net reaction :Zno(s)+Fe 2+ = Zn 3+ Feo(s)
1
Expert's answer
2020-06-22T08:52:15-0400

oxidation Zn(s) → Zn3+(aq) + 2e–

E° = 0.763 V

reduction

Fe2+(aq)+2e- → Fe(s)=-0.44V

Ecell=Eoxd-Ered=0.763-(-0.44)=1.203V